Determine whether the following force field
Fis conservative:
F=xi-yj+zk
1
Expert's answer
2020-03-09T10:55:54-0400

The conservative field "\bf F" must satisfy condition


"\rm rot\:{\bf F}=0 ."

We have


"{\rm rot}\:(x\hat i-y\hat j+z\hat k)=\begin{vmatrix}
 \hat i & \hat j & \hat k \\
 x & -y & z\\
\frac{\partial}{\partial x} & \frac{\partial}{\partial y}& \frac{\partial}{\partial z}
\end{vmatrix}"

"=\hat i\left(\frac{\partial (-y)}{\partial z}-\frac{\partial z}{\partial y})\right)-\hat j\left(\frac{\partial x}{\partial z}-\frac{\partial z}{\partial x})\right)+"

"+\hat k\left(\frac{\partial x}{\partial y}-\frac{\partial (-y)}{\partial x})\right)=0\hat i+0\hat j+0\hat k=\bf 0."

Hence, the field "\bf F" is conservative.